HÍRDETÉSEK

1. Évfolyam ZH

Az évfolyam ZH időpontjai:

1. 2006. május 8-án 17h-tól 20h-ig,

2. 2006. május 9-én 17h-tól 20h-ig,

3. 2006. május 11-én 17h-tól 20h-ig.

4. 2006. május 12-én 17h-tól 20h-ig.

Az évfolyam ZH helye: LOVARDA. Budapest, 2006. február 13.

Tőke Pál sk.